A random variable X has the probability density function given below

f(x)="\begin{Bmatrix}
 x, 0<=x<1 \\
 2-x, 1<=x<2\\
0, otherwise
\end{Bmatrix}" Show graphically and find out the robabilities. p(-1<x<55)


1
Expert's answer
2021-08-16T14:11:17-0400

"f(x) = \begin{cases}
 0 &x\leq 0\\
 x &0\leq x< 1 \\
 2-x &1\leq x< 2 \\
 0 &x\geq 2 \\
\end{cases}"

"\displaystyle\int_{-\infin}^{\infin}f(y)dy=\dfrac{1}{2}(2-0)(1)=1"



"F(x)=\displaystyle\int_{-\infin}^{x}f(y)dy"

"P(a\leq X\leq b)=F(b)-F(a)"

Then


"P(-1<X<55)=1"